### v3

# Structural Asymmetry and Rigidity of Bounded Solutions for a Parametrized Complex Linear Differential Equation

Abstract
We introduce a family of parametrized non-homogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter within the critical strip. We establish a rigidity phenomenon: Two solutions with the same initial condition equal to 11, corresponding to two different parameters lying on a horizontal line in the critical strip, cannot both be bounded on $1,+\infty)$. In particular, when $\Re(w)\neq\tfrac{1}{2}$, the two solutions corresponding respectively to the parameters $w$ and $1-\bar{w}$, both taken with initial condition 11, cannot be simultaneously bounded. This result provides an answer to the Dynamical Conjecture formulated in [2].
Keywords: Non-homogeneous complex linear differential equation, bounded solutions.
MSC subject classifications: 34A30, 34E05
1 Main results
In this manuscript, we study the following non-homogeneous linear complex parametrized differential equation
$\dot{x}=w\,t^{-1}x+(1-w)\,t^{-1}\eta(t),$ (1)
$t\in[1,+\infty),\quad x(1)=z,\quad x:[1,+\infty)\to\mathbb{C},$
where the parameter $w\in\mathbb{C}$. The the nonhomogeneous term $\eta:[1,+\infty)\to\mathbb{C}$ is a locally integrable function satisfying the following hypothesis (H):
$|\eta(t)|<1,\quad\forall t\geq 1,\quad\text{and}\quad\exists\rho\in\mathbb{C}^{*}:\ \sup_{t\geq 1}\left|\int_{1}^{t}(\eta(u)-\rho)\,du\right|<+\infty.$ (H)
We call $\rho$ the rotation number of the function $\eta$. The equation (1) can be viewed as a differential equation defined on $1,+\infty)$ with values in $\mathbb{R}^{2}$, and we are interested in the initial conditions of the bounded solutions. More precisely, our aim is to prove that for every $w,w^{\prime}\in\mathbb{C}$ with $\Re(w)\neq\Re(w)$ and $\Im(w)=\Im(w^{\prime})$, the solution $x_{w}$ and the solution $x_{w^{\prime}}$ with initial conditions $x_{w}(1)=x_{1-\overline{w}}(1)=1$ cannot both be bounded on $1,+\infty)$. In the study that follows, we fix a specific function $\eta$ which is locally integrable satisfying hypothesis (H) with a fixed rotation number $\rho\in\mathbb{C}^{*}$. The function
$t\mapsto t^{w}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1,$
is continuous on $1,+\infty)$. The differential equation (1) is a non-homogeneous linear differential equation. Then, for every $z\in\mathbb{C}$, there exists a unique continuous solution $\psi_{w}(z,\cdot)$ of (1) such that $\psi_{w}(z,1)=z$, which is given by
$\psi_{w}(z,t)=t^{w}\Big[z+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big],\quad\forall t\geq 1.$ (2)
The analysis in this manuscript is valid only when the real part of the parameter is in $(0,1)$. We consider the following parameter set,
$B:=\Big\{w\in\mathbb{C}:\quad\Re(w)\in(0,1),\quad\Im(w)\neq 0\Big\}.$
Let us introduce some notations.
Notation 1.
For every $w\in B$ and $z\in\mathbb{C}$, we denote by $\psi_{w}(z,\cdot):[1,+\infty)\to\mathbb{C}$ the unique continuous solution of the differential equation (1). The point $z$ is called the initial condition of the solution $\psi_{w}(z,\cdot)$.
Notation 2.
We denote by $\mu$ the function $\mu:B\mapsto\mathbb{C}$ defined as
$\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\eta(u)\,du,\quad\forall s\in B.$
In the following lemma, we are interested in bounded solutions on $1,+\infty)$ and their initial condition.
Lemma 3.
Let $w\in B$ and $z\in\mathbb{C}$. Consider the continuous solution $\psi_{w}(z,\cdot)$ of the differential equation (1) as given in Notation 1. Then
$\sup_{t\geq 1}|\psi_{w}(z,t)|<+\infty\quad\text{if and only if}\quad z=\mu(w),$
where the function $\mu$ is given by the Notation 2. Further,
$\sup_{t\geq 1}\left|t\Big(\psi_{{w}}(z_{{w}},t)+\rho\dfrac{1-w}{w}\Big)\right|<+\infty.$
Proof.
Let $w\in B$ be fixed. There is a unique bounded solution; all other solutions are oscillating and diverge in norm. Let $z\in\mathbb{C}$ and suppose that $\sup_{t\geq 1}|\psi_{w}(z,t)|<+\infty$. By equation (2), this implies
$\sup_{t\geq 1}\Bigg[t^{\Re(w)}\Big|z+(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du\Big|\Bigg]<+\infty.$
Since $\Re(w)>0$, we must have
$\Big|z+(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du\Big|=0,$
which gives $z=\mu(w)$. To prove that $\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|<+\infty$, observe that by definition of $\mu(w)$,
$\mu(w)+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du=-(1-w)\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$
Using equation (2), we then obtain
$\psi_{w}(\mu(w),t)=-(1-w)\,t^{w}\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$ (3)
From the Hypothesis (H) we have $\sup_{t\geq 1}|\eta(t)|\leq 1$, we obtain
$\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|\leq\frac{|1-w|}{\Re(w)}\sup_{t\geq 1}|\eta(t)|<\frac{|1-w|}{\Re(w)}.$
Now, prove the last item of the present Lemma. From equation (3), we have
$\psi_{{w}}(z_{{w}},t)=-\rho\dfrac{1-w}{w}-(1-w)t^{{w}}\int_{t}^{+\infty}u^{-1-w}(\eta(u)-\rho)du,\quad\forall t\geq 1.$
Using the integration by parts formula, we obtain
$\psi_{{w}}(z_{{w}},t)=-\rho\dfrac{1-w}{w}-(1-w^{2})t^{w}\int_{t}^{+\infty}u^{-2-w}\int_{1}^{u}(\eta(v)-\rho)dvdu.$
For every $t\geq 1$, we get
$\left|\psi_{{w}}(z_{{w}},t)+\rho\dfrac{1-w}{w}\right|\leq\,t^{-1}\dfrac{|1-w^{2}|}{1+\Re(w)}\sup_{t\geq u}\Big|\int_{1}^{u}(\eta(v)-\rho)dv\Big|.$
By the Hypothesis (H), we obtain
$\sup_{t\geq 1}\left|t\Big(\psi_{{w}}(z_{{w}},t)+\rho\dfrac{1-w}{w}\Big)\right|<+\infty.$
The following theorem, show in particular, that if $\Re(s)\neq\frac{1}{2}$, then the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) cannot both be bounded on $1,+\infty)$.
Theorem 4.
Consider the function $\mu$ given by the Notation 2. For every $s,s^{\prime}\in B$ with $\Re(s)\neq\Re(s^{\prime})$ and $\Im(s)=\Im(s^{\prime})$, we have
$\big(\mu(s),\,\mu(s^{\prime}))\neq(1,1).$
In particular, for every $s\in B$ we have the following implication:
$\big(\mu(s),\,\mu(1-\overline{s})\big)=(1,1)\implies\Re(s)=\frac{1}{2}.$
2 Main Proposition
From the perspective of the qualitative theory of dynamical systems, the differential equation (1) can be viewed as a differential equation with the parameter $\Re(w)$. Intuitively, this introduces a order structure within the set of solutions with a same initial condition and parameterized by $\Re(w)$. We show this in the following proposition.
Proposition 5.
Let be $s:=\sigma+i\tau\in B$ and $s^{\prime}:=\sigma^{\prime}+i\tau\in B$. Consider the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{s^{\prime}}(1,\cdot)$ of the differential equation (1) as given in Notation 1. Suppose that $\sigma>\sigma^{\prime}$, then
$\Re\big(t^{-i\,\tau}\psi_{s}(1,t)\big)>\Re\big(t^{-i\,\tau}\psi_{s^{\prime}}(1,t)\big),\qquad\forall t>1.$
Proof.
Let be $w\in B$ and denote $w:=\alpha+i\beta$ . For every $t\geq 1$ define
$\psi_{w}(t):=t^{-i\beta}\psi_{w}(1,t).$
According to Notation 1, these functions satisfy the differential equations
$\dot{\psi}_{w}(t)=\alpha\,t^{-1}\psi_{w}(t)+(1-w)\,t^{-1-i\beta}\eta(t),\quad\psi_{w}(1)=1.$
In order to simplify the notation, denote
$x_{w}(t):=\Re\big(\psi_{w}(t)\big)-1,\qquad\forall t\geq 1.$
The previous differential equations, implies that
$\dot{x}_{w}(t)=\alpha\,t^{-1}x_{w}(t)+\alpha t^{-1}+t^{-1}\Re\big((1-w)\,t^{-i\beta}\,\eta(t)\big),\quad x_{w}(1)=0.$ (4)
Let be $s:=\sigma+i\tau\in B$ and $s:=\sigma^{\prime}+i\tau\in B$ such that $\sigma>\sigma^{\prime}$. The previous differential equation implies
$\dot{x}_{s}(t)-\dot{x}_{s^{\prime}}(t)$ $=\sigma\,t^{-1}x_{s}(t)-\sigma^{\prime}\,t^{-1}x_{s^{\prime}}(t)$ (5)
$+(\sigma-\sigma^{\prime})t^{-1}\big(1-\Re\big(\,t^{-i\tau}\,\eta(t)\big)\big).$
We have $x_{s}(1)=x_{s^{\prime}}(1)=0$. The previous differential equation implies that
$\dot{x}_{s}(1)-\dot{x}_{s^{\prime}}(1)=(\sigma-\sigma^{\prime})t^{-1}\big(1-\Re\big(\eta(t)\big)\big).$
By the hypothesis (H), we have $|\eta(t)|<1$ for all $t\geq 1$. By hypothesis we have $\sigma>\sigma^{\prime}$, then $\dot{x}_{s}(1)>\dot{x}_{s^{\prime}}(1)$. By continuity there exists $\epsilon>0$ such that
$x_{s}(t)>x_{s^{\prime}}(t),\quad\forall t\in(1,1+\epsilon).$
Define the time $t_{*}>1$ as
$t_{*}:=\sup\{t>1:\forall\ 1<u<t,\ x_{s}(u)>x_{s^{\prime}}(u)\}.$ (6)
The strategy to prove the present Proposition is to prove that $t_{*}=+\infty$. By contradiction, suppose that $t_{*}<+\infty$. Then
$x_{s}(t_{*})=x_{s^{\prime}}(t_{*}),$
and we have
$x_{s}(t)>x_{s^{\prime}}(t),\quad\forall t\in(1,t_{*}).$ (7)
We have two cases: $x_{s}(t_{*})\geq 0$ and $x_{s}(t_{*})<0$.
If $x_{s}(t_{*})\geq 0$, then $x_{s^{\prime}}(t_{*})\geq 0$. Using the fact that $\sigma>\sigma^{\prime}$, from the differential equations (5), we get
$\dot{x}_{s}(t_{*})-\dot{x}_{s^{\prime}}(t_{*})$ $=(\sigma-\sigma^{\prime})t_{*}^{-1}x_{s}(t_{*})$
$+(\sigma-\sigma^{\prime})t_{*}^{-1}\big(1-\Re\big(\,t_{*}^{-i\tau}\,\eta(t_{*})\big)\big)$
$\geq(\sigma-\sigma^{\prime})t_{*}^{-1}\big(1-\Re\big(\,t_{*}^{-i\tau}\,\eta(t_{*})\big)\big).$
By the hypothesis (H), we obtain $\dot{x}_{s}(t_{*})>\dot{x}_{s^{\prime}}(t_{*})$. There exists $t<t_{*}$ close enough to $t_{*}$ such that $x_{s}(t)<x_{s^{\prime}}(t)$. We have obtained a contradiction with equation (7).
If $x_{s}(t_{*})<0$, then $x_{s^{\prime}}(t_{*})<0$. By continuity there exists $\tilde{\epsilon}>0$ such that
$0>x_{s}(u)>x_{s^{\prime}}(u),\quad\forall u\in(t_{*}-\tilde{\epsilon},t_{*}).$
Implies
$\delta_{s,s^{\prime}}(u):=x_{s}(u)x_{s^{\prime}}(u)-(x_{s}(u))^{2}=x_{s}(u)(x_{s^{\prime}}(u)-x_{s}(u))>0,\ \forall u\in(t_{*}-\tilde{\epsilon},t_{*}).$
and $\delta_{s,s^{\prime}}(t_{*})=0$. Then we must find $\dot{\delta}_{s,s^{\prime}}(t_{*})\leq 0$. As the previous technique, prove that in our system that $\dot{\delta}_{s,s^{\prime}}(t_{*})>0$. From the differential equation (4), we have
[…]
$x_{s^{\prime}}(t)\dot{x}_{s}(t)$ $=\sigma\,t^{-1}x_{s^{\prime}}(t)x_{s}(t)+\sigma t^{-1}x_{s^{\prime}}(t)+t^{-1}x_{s^{\prime}}(t)\Re\big((1-s)\,t^{-i\tau}\,\eta(t)\big),$
$x_{s}(t)\dot{x}_{s^{\prime}}(t)$ $=\sigma^{\prime}\,t^{-1}x_{s}(t)x_{s^{\prime}}(t)+\sigma^{\prime}t^{-1}x_{s}(t)+t^{-1}x_{s}(t)\Re\big((1-s^{\prime})\,t^{-i\tau}\,\eta(t)\big).$
Implies
$\dot{\delta}_{s,s^{\prime}}(t)$ $=2\sigma\,t^{-1}{\delta}_{s,s^{\prime}}(t)+t^{-1}\big(\sigma x_{s^{\prime}}(t)+\sigma^{\prime}x_{s}(t)\big)$
$+t^{-1}x_{s^{\prime}}(t)\Re\big((1-s)\,t^{-i\tau}\,\eta(t)\big)$
$-t^{-1}x_{s}(t)\Re\big((1-s^{\prime})\,t^{-i\tau}\,\eta(t)\big)$
$-2\sigma t^{-1}x_{s}(t).$
At the point $t_{*}$ we get
$\dot{\delta}_{s,s^{\prime}}(t_{*})$ $=-t_{*}^{-1}\big(\sigma-\sigma^{\prime}\big)x_{s}(t_{*})\big(1-\Re\big(t_{*}^{-i\tau}\,\eta(t_{*})\big)\big).$
Use the fact that $\sigma>\sigma^{\prime}$ and $x_{s}(t_{*})<0$, we obtain $\dot{\delta}_{s,s^{\prime}}(t_{*})>0$. Same, we obtained a contradiction. Then $t_{*}=+\infty$. ∎
3 Proof of the Theorem 4
Intuitively, the solution $\psi_{w}(1,\cdot)$cannot be bounded when $\Re(w)\in\left(\tfrac{1}{2},1\right)$. Indeed, if $\Re(w)\in(0,1)$, then |w|>|1w||w|>|1-w|. Since the initial condition satisfies $\psi_{w}(1,1)=1$ and $\sup_{t\geq 1}|\eta(t)|\leq 1$, for $t$ close to 11 the homogeneous part dominates the non-homogeneous part in differential equation (1). This may therefore lead to the unboundedness of the solution $\psi_{w}(1,\cdot)$. In this manuscript, we compare and study the boundedness of two solutions corresponding to two complex parameters with the same imaginary part, in order to eliminate the imaginary part (which may be very large) from the computations.
Proof of the Theorem 4.
We prove the theorem by contradiction. Suppose there exist $s,s^{\prime}\in B$ with $\Re(s)>\Re(s^{\prime})$ and $\Im(s)=\Im(s^{\prime})$ such that $\big(\mu(s),\,\mu(s^{\prime})\big)=(1,1)$. The solutions $\psi_{s}(1,\cdot)$ and $\psi_{s^{\prime}}(1,\cdot)$ of the differential equation (1) are both bounded on $1,+\infty)$. Thanks to the second Item of the Lemma 3, we have
$\sup_{t\geq 1}\Big|t\Big(\psi_{s}(1,t)+\rho\frac{1-s}{s}\Big)\Big|<+\infty\ \text{and}\ \sup_{t\geq 1}\Big|t\Big(\psi_{s^{\prime}}(1,t)+\rho\frac{1-s^{\prime}}{s^{\prime}}\Big)\Big|<+\infty.$
Use the Proposition 5, we get
$\lim_{t\to+\infty}\Re\Big(t^{-i\Im(s)}\rho\big(\frac{1-s^{\prime}}{s^{\prime}}-\frac{1-s}{s}\big)\Big)\geq 0.$
Since $s,s^{\prime}\in B$, then $\Im(s)\neq 0$, that gives
$\rho\big(\frac{1-s^{\prime}}{s^{\prime}}-\frac{1-s}{s}\big)=0.$
By the hypothesis (H), we have $\rho\neq 0$. then
$\frac{1-s^{\prime}}{s^{\prime}}-\frac{1-s}{s}=0$
We have obtained a contradiction with $\Re(s)\neq\Re(s^{\prime})$. ∎
4 Application to the $\zeta$-Function
The function $\eta(t)=\{t\}$ (the fractional part of $t$) is a locally integrable function and satisfies $0\leq\eta(t)<1$ with a rotation number $\rho$ (here $\rho=1/2$). The Hypothesis (H) is satisfied. By definition of the function $\mu$ in the Definition 2, we have
$\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\{u\}\,du,\quad\mu(1-\bar{s})=-\bar{s}\int_{1}^{+\infty}u^{-2+\bar{s}}\{u\}\,du.$
In other words, by considering the integral representation of the Riemann zeta function $\zeta$ defined by Abel’s summation formula [[1], page 14, Equation 2.1.5], we have the relations
$\frac{1-s}{s}\zeta(s):=-1+\mu(s),\quad\frac{\bar{s}}{1-\bar{s}}\zeta(1-\bar{s}):=-1+\mu(1-\bar{s}).$ (8)
Thanks to the Riemann functional equation [1, Theorem 2.1, p. 13], when $\zeta(s)=0$ then $\zeta(1-\bar{s})=0$. Implies $\big(\mu(s),\ \mu(1-\bar{s})\big)=(1,1)$. By the Theorem 4 we deduce that $\Re(s)=\Re(1-\bar{s})$. The non-trivial zeros of the Riemann zeta function must lie on the critical line.

### v4

@@
 Abstract
-We introduce a family of parametrized non-homogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter within the critical strip.
-We establish a rigidity phenomenon: Two solutions with the same initial condition equal to 11, corresponding to two different parameters lying on a horizontal line in the critical strip, cannot both be bounded on $1,+\infty)$.
-In particular, when $\Re(w)\neq\tfrac{1}{2}$, the two solutions corresponding respectively to the parameters $w$ and $1-\bar{w}$, both taken with initial condition 11, cannot be simultaneously bounded.
-This result provides an answer to the Dynamical Conjecture formulated in [2].
+In this manuscript, we introduce a family of parametrized nonhomogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter lying in the critical strip.
+We identify a suitable rigidity hypothesis under which a phenomenon of structural asymmetry is established: if two solutions with the same initial condition equal to 11, corresponding respectively to the parameters $s$ and $1-\bar{s}$, are both bounded on $1,+\infty)$, then $\Re(s)=\frac{1}{2}$.
 Keywords: Non-homogeneous complex linear differential equation, bounded solutions.
@@
 1 Main results
-In this manuscript, we study the following non-homogeneous linear complex parametrized differential equation
-$\dot{x}=w\,t^{-1}x+(1-w)\,t^{-1}\eta(t),$ (1)
+In this manuscript, we study the following non-homogeneous linear complex parametrized differential equations
+$\dot{x}=\Re(w)\,t^{-1}x+(1-w)\,t^{-1-i\,\Im(w)}\eta(t),$ (1)
 $t\in[1,+\infty),\quad x(1)=z,\quad x:[1,+\infty)\to\mathbb{C},$
-where the parameter $w\in\mathbb{C}$. The the nonhomogeneous term $\eta:[1,+\infty)\to\mathbb{C}$ is a locally integrable function satisfying the following hypothesis (H):
-$|\eta(t)|<1,\quad\forall t\geq 1,\quad\text{and}\quad\exists\rho\in\mathbb{C}^{*}:\ \sup_{t\geq 1}\left|\int_{1}^{t}(\eta(u)-\rho)\,du\right|<+\infty.$ (H)
-We call $\rho$ the rotation number of the function $\eta$.
+where $w\in\mathbb{C}$ and $\eta:[1,+\infty)\to\mathbb{C}$ is a locally integrable function and satisfy $|\eta(t)|<1$ for all $t\geq 1$.
 The equation (1) can be viewed as a differential equation defined on $1,+\infty)$ with values in $\mathbb{R}^{2}$, and we are interested in the initial conditions of the bounded solutions.
-More precisely, our aim is to prove that for every $w,w^{\prime}\in\mathbb{C}$ with $\Re(w)\neq\Re(w)$ and $\Im(w)=\Im(w^{\prime})$, the solution $x_{w}$ and the solution $x_{w^{\prime}}$ with initial conditions $x_{w}(1)=x_{1-\overline{w}}(1)=1$ cannot both be bounded on $1,+\infty)$.
-In the study that follows, we fix a specific function $\eta$ which is locally integrable satisfying hypothesis (H) with a fixed rotation number $\rho\in\mathbb{C}^{*}$.
+This result provides an answer to the Dynamical Conjecture formulated in [2].
+In the study that follows, we fix a specific function $\eta$ which is locally integrable function and satisfy $|\eta(t)|<1$ for all $t\geq 1$.
 The function
-$t\mapsto t^{w}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1,$
+$t\mapsto t^{\Re(w)}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1,$
 is continuous on $1,+\infty)$.
@@
 Then, for every $z\in\mathbb{C}$, there exists a unique continuous solution $\psi_{w}(z,\cdot)$ of (1) such that $\psi_{w}(z,1)=z$, which is given by
-$\psi_{w}(z,t)=t^{w}\Big[z+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big],\quad\forall t\geq 1.$ (2)
-The analysis in this manuscript is valid only when the real part of the parameter is in $(0,1)$. We consider the following parameter set,
-$B:=\Big\{w\in\mathbb{C}:\quad\Re(w)\in(0,1),\quad\Im(w)\neq 0\Big\}.$
+$\psi_{w}(z,t)=t^{\Re(w)}\Big[z+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big],\quad\forall t\geq 1.$ (2)
 Let us introduce some notations.
 Notation 1.
-For every $w\in B$ and $z\in\mathbb{C}$, we denote by $\psi_{w}(z,\cdot):[1,+\infty)\to\mathbb{C}$ the unique continuous solution of the differential equation (1).
+For every $w\in\mathbb{C}$ and $z\in\mathbb{C}$, we denote by $\psi_{w}(z,\cdot):[1,+\infty)\to\mathbb{C}$ the unique continuous solution of the differential equation (1) given by equation (2).
 The point $z$ is called the initial condition of the solution $\psi_{w}(z,\cdot)$.
 Notation 2.
-We denote by $\mu$ the function $\mu:B\mapsto\mathbb{C}$ defined as
+Denote by $B\subset\mathbb{C}$ the following set
+$B:=\Big\{w\in\mathbb{C}:\quad\Re(w)>0\Big\}.$
+Let $\mu$ denote the function $\mu:B\mapsto\mathbb{C}$, defined as
 $\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\eta(u)\,du,\quad\forall s\in B.$
@@
 $\sup_{t\geq 1}|\psi_{w}(z,t)|<+\infty\quad\text{if and only if}\quad z=\mu(w),$
-where the function $\mu$ is given by the Notation 2. Further,
-$\sup_{t\geq 1}\left|t\Big(\psi_{{w}}(z_{{w}},t)+\rho\dfrac{1-w}{w}\Big)\right|<+\infty.$
+where the function $\mu$ is given by the Notation 2.
 Proof.
 Let $w\in B$ be fixed.
-There is a unique bounded solution; all other solutions are oscillating and diverge in norm.
+There is a unique bounded solution; all other solutions diverge in norm.
 Let $z\in\mathbb{C}$ and suppose that $\sup_{t\geq 1}|\psi_{w}(z,t)|<+\infty$.
 By equation (2), this implies
-$\sup_{t\geq 1}\Bigg[t^{\Re(w)}\Big|z+(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du\Big|\Bigg]<+\infty.$
+$\sup_{t\geq 1}\Bigg[t^{\Re(w)}\Big|z+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big|\Bigg]<+\infty.$
 Since $\Re(w)>0$, we must have
@@
 Using equation (2), we then obtain
-$\psi_{w}(\mu(w),t)=-(1-w)\,t^{w}\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$ (3)
-From the Hypothesis (H) we have $\sup_{t\geq 1}|\eta(t)|\leq 1$, we obtain
+$\psi_{w}(\mu(w),t)=-(1-w)\,t^{\Re(w)}\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$ (3)
+By Hypothesis we have $\sup_{t\geq 1}|\eta(t)|\leq 1$, we obtain
 $\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|\leq\frac{|1-w|}{\Re(w)}\sup_{t\geq 1}|\eta(t)|<\frac{|1-w|}{\Re(w)}.$
-Now, prove the last item of the present Lemma. From equation (3), we have
-$\psi_{{w}}(z_{{w}},t)=-\rho\dfrac{1-w}{w}-(1-w)t^{{w}}\int_{t}^{+\infty}u^{-1-w}(\eta(u)-\rho)du,\quad\forall t\geq 1.$
-Using the integration by parts formula, we obtain
-$\psi_{{w}}(z_{{w}},t)=-\rho\dfrac{1-w}{w}-(1-w^{2})t^{w}\int_{t}^{+\infty}u^{-2-w}\int_{1}^{u}(\eta(v)-\rho)dvdu.$
-For every $t\geq 1$, we get
-$\left|\psi_{{w}}(z_{{w}},t)+\rho\dfrac{1-w}{w}\right|\leq\,t^{-1}\dfrac{|1-w^{2}|}{1+\Re(w)}\sup_{t\geq u}\Big|\int_{1}^{u}(\eta(v)-\rho)dv\Big|.$
-By the Hypothesis (H), we obtain
-$\sup_{t\geq 1}\left|t\Big(\psi_{{w}}(z_{{w}},t)+\rho\dfrac{1-w}{w}\Big)\right|<+\infty.$
-The following theorem, show in particular, that if $\Re(s)\neq\frac{1}{2}$, then the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) cannot both be bounded on $1,+\infty)$.
-Theorem 4.
-Consider the function $\mu$ given by the Notation 2. For every $s,s^{\prime}\in B$ with $\Re(s)\neq\Re(s^{\prime})$ and $\Im(s)=\Im(s^{\prime})$, we have
-$\big(\mu(s),\,\mu(s^{\prime}))\neq(1,1).$
-In particular, for every $s\in B$ we have the following implication:
-$\big(\mu(s),\,\mu(1-\overline{s})\big)=(1,1)\implies\Re(s)=\frac{1}{2}.$
-2 Main Proposition
 From the perspective of the qualitative theory of dynamical systems, the differential equation (1) can be viewed as a differential equation with the parameter $\Re(w)$.
@@
 Proposition 4.
-Let be $s:=\sigma+i\tau\in B$ and $s^{\prime}:=\sigma^{\prime}+i\tau\in B$.
-Consider the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{s^{\prime}}(1,\cdot)$ of the differential equation (1) as given in Notation 1.
-Suppose that $\sigma>\sigma^{\prime}$, then
-$\Re\big(t^{-i\,\tau}\psi_{s}(1,t)\big)>\Re\big(t^{-i\,\tau}\psi_{s^{\prime}}(1,t)\big),\qquad\forall t>1.$
-Proof.
-Let be $w\in B$ and denote $w:=\alpha+i\beta$ . For every $t\geq 1$ define
-$\psi_{w}(t):=t^{-i\beta}\psi_{w}(1,t).$
-According to Notation 1, these functions satisfy the differential equations
-$\dot{\psi}_{w}(t)=\alpha\,t^{-1}\psi_{w}(t)+(1-w)\,t^{-1-i\beta}\eta(t),\quad\psi_{w}(1)=1.$
-In order to simplify the notation, denote
-$x_{w}(t):=\Re\big(\psi_{w}(t)\big)-1,\qquad\forall t\geq 1.$
-The previous differential equations, implies that
-$\dot{x}_{w}(t)=\alpha\,t^{-1}x_{w}(t)+\alpha t^{-1}+t^{-1}\Re\big((1-w)\,t^{-i\beta}\,\eta(t)\big),\quad x_{w}(1)=0.$ (4)
-Let be $s:=\sigma+i\tau\in B$ and $s:=\sigma^{\prime}+i\tau\in B$ such that $\sigma>\sigma^{\prime}$. The previous differential equation implies
-$\dot{x}_{s}(t)-\dot{x}_{s^{\prime}}(t)$ $=\sigma\,t^{-1}x_{s}(t)-\sigma^{\prime}\,t^{-1}x_{s^{\prime}}(t)$ (5)
-$+(\sigma-\sigma^{\prime})t^{-1}\big(1-\Re\big(\,t^{-i\tau}\,\eta(t)\big)\big).$
-We have $x_{s}(1)=x_{s^{\prime}}(1)=0$. The previous differential equation implies that
-$\dot{x}_{s}(1)-\dot{x}_{s^{\prime}}(1)=(\sigma-\sigma^{\prime})t^{-1}\big(1-\Re\big(\eta(t)\big)\big).$
-By the hypothesis (H), we have $|\eta(t)|<1$ for all $t\geq 1$.
-By hypothesis we have $\sigma>\sigma^{\prime}$, then $\dot{x}_{s}(1)>\dot{x}_{s^{\prime}}(1)$.
-By continuity there exists $\epsilon>0$ such that
-$x_{s}(t)>x_{s^{\prime}}(t),\quad\forall t\in(1,1+\epsilon).$
+Let be $s:\sigma+i\tau\in B$.
+Consider the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) as given in Notation 1.
+The function $\delta_{s}$ defined as
+$\delta_{s}(t):=\Re\big({\psi_{s}}(1,t)-{\psi}_{1-\bar{s}}(1,t)\big),\quad\forall t\geq 1.$
+satisfies the following two equalities:
+$\dot{\delta}_{s}(1)$ $=(2\sigma-1)\Big(1-\Re\big(\eta(1)\big)\Big),$
+${\delta}_{s}(t)$ $=\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
+$\ +4(\sigma-\frac{1}{2})\Re\!\left(\psi_{\frac{1}{2}+i\,\tau}(1,t)-\psi_{i\tau}(1,t)\right),\qquad\forall t\geq 1.$
+According to the notation 1 consider the following rigidity hypothesis H:
+  • H.
+    For every $\beta\in\mathbb{R}$ and every $t>1$,
+    $\Re\!\left(\psi_{\frac{1}{2}+i\beta}(1,t)-\psi_{i\beta}(1,t)\right)>0.$
+The previous hypothesis suggest the rigidity of the sign. It is suitable in order to apply the results of this dynamical study to the Riemann zeta function.
+The following theorem shows that, under the rigidity hypothesis H, if $\Re(s)\neq\tfrac{1}{2}$, then the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) cannot both be bounded on $1,+\infty)$.
+Theorem 5.
+Consider the function $\mu$ defined in Notation 2. Under hypothesis H, for every $s\in\mathbb{C}$ with $\Re(s)\in(0,1)/\{\frac{1}{2}\}$, we have
+$\big(\mu(s),\,\mu(1-\overline{s})\big)\neq(1,1).$
+2 Proof of the Proposition 4 and Theorem 5
+Proof of the Proposition 4.
+Let be $w\in B$ and denote $w:=\alpha+i\beta$. For every $t\geq 1$ denote
+$x_{w}(t):=t^{-\frac{1}{2}}\Re\big(\psi_{w}(1,t)\big).$ (4)
+According to Notation 1, we obtain the following differential equation
+$\dot{x}_{w}(t)=\big(\alpha-\frac{1}{2}\big)\,t^{-1}x_{w}(t)+t^{-\frac{3}{2}}\Re\big((1-w)\,t^{-i\beta}\,\eta(t)\big),\quad x_{w}(1)=1.$ (5)
+Let be $s:=\sigma+i\tau\in B$ such that $\sigma\in(\frac{1}{2},1)$. From the previous equation, we have
+$\dot{x_{s}}(t)+\dot{x}_{1-\bar{s}}(t)$ $=(\sigma-\frac{1}{2})t^{-1}\big({x_{s}}(t)-{x}_{1-\bar{s}}(t)\big)+t^{-\frac{3}{2}}\Re\Big((1-2i\tau)t^{-i\tau}\,\eta(t)\Big).$
+Use the fact that $x_{s}(1)=x_{1-\bar{s}}(1)=1$ and integrate, we obtain
+${x_{s}}(t)+{x}_{1-\bar{s}}(t)$ $=(\sigma-\frac{1}{2})\int_{1}^{t}u^{-1}\big({x_{s}}(u)-{x}_{1-\bar{s}}(u)\big)\,du$
+$+2+\Re\Big((1-2i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\tau}\,\eta(u)\,du\Big).$
+From the differential equation (5), we have
+$\dot{x}_{s}(t)-\dot{x}_{1-\bar{s}}(t)$ $=(\sigma-\frac{1}{2})t^{-1}\big(x_{s}(t)+x_{1-\bar{s}}(t)\big)$
+$-(2\sigma-1)t^{-\frac{3}{2}}\Re\big(\,t^{-i\tau}\,\eta(t)\big).$
+The two previous differential equations, gives
+$\dot{x}_{s}(t)-\dot{x}_{1-\bar{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{-1}\int_{1}^{t}u^{-1}\big({x_{s}}(u)-{x}_{1-\bar{s}}(u)\big)\,du$
+$+2(\sigma-\frac{1}{2})t^{-1}\Big(1-t^{-\frac{1}{2}}\Re\big(\,t^{-i\tau}\,\eta(t)\big)\Big)$
+$+(\sigma-\frac{1}{2})t^{-1}\Re\Big((1-2i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\tau}\,\eta(u)\,du\Big).$
+Define
+$\delta_{s}(t):=\Re\big({\psi_{s}}(1,t)-{\psi}_{1-\bar{s}}(1,t)\big),\quad\forall t\geq 1.$
+By the notation (4), we have ${x_{s}}(t)-{x}_{1-\bar{s}}(t):=t^{-\frac{1}{2}}\delta_{s}(t)$. We obtained
+$\dot{\delta}_{s}(t)=$ $\frac{1}{2}t^{-1}\delta_{s}(t)+(\sigma-\frac{1}{2})^{2}t^{-\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}}\delta_{s}(u)\,du$
+$+2(\sigma-\frac{1}{2})t^{-\frac{1}{2}}\Big(1-t^{-\frac{1}{2}}\Re\big(\,t^{-i\tau}\,\eta(t)\big)\Big)$
+$+2(\sigma-\frac{1}{2})t^{-\frac{1}{2}}\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\tau}\,\eta(u)\,du\Big).$
+Since ${\delta}_{s}(1)=0$ and since by hypothesis we have $|\eta(1)|<1$, then
+$\dot{\delta}_{s}(1)=(2\sigma-1)\Big(1-\Re\big(\eta(1)\big)\Big).$
+Now, prove the second item of the present Proposition. Integrate the previous equation, we obtain
+${\delta}_{s}(t)=\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}\dot{\delta}_{s}(v)\,dv\,du$
+$+2(\sigma-\frac{1}{2})\int_{1}^{t}u^{-\frac{1}{2}}\Re\Big(1-u^{-\frac{1}{2}-i\tau}\,\eta(u)\Big)du$
+$+2(\sigma-\frac{1}{2})\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}-i\tau}\,\eta(v)\,dv\,du\Big).$
+Use the integration by pars formula to the last integral, we get
+${\delta}_{s}(t)=\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}\dot{\delta}_{s}(v)\,dv\,du$
+$+2(\sigma-\frac{1}{2})\int_{1}^{t}u^{-\frac{1}{2}}\Re\Big(1-u^{-\frac{1}{2}-i\tau}\,\eta(u)\Big)du$
+$+4(\sigma-\frac{1}{2})t^{\frac{1}{2}}\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}v^{-\frac{3}{2}-i\tau}\,\eta(v)\,dv\Big)$
+$-4(\sigma-\frac{1}{2})\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-1-i\tau}\,\eta(u)\,du\Big).$
+Equivalent to
+${\delta}_{s}(t)=$ $\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
+$+4(\sigma-\frac{1}{2})t^{\frac{1}{2}}\big[(\frac{1}{2}-i\tau)\int_{1}^{t}v^{-\frac{3}{2}-i\tau}\,\eta(v)\,dv\Big]$
+$-4(\sigma-\frac{1}{2})\Big[1+\Re\Big((1-i\tau)\int_{1}^{t}u^{-1-i\tau}\,\eta(u)\,du\Big)\Big].$
+According to Notation 1, we obtain
+${\delta}_{s}(t)=$ $\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
+$+4(\sigma-\frac{1}{2})\Re\!\left(\psi_{\frac{1}{2}+i\beta}(1,t)-\psi_{i\beta}(1,t)\right).$
+Proof of the Theorem 5.
+Let be $s:=\sigma+i\tau\in B$ such that $\sigma\in(0,1)/\{\frac{1}{2}\}$. By symmetry, we may assume that $\sigma\in(\frac{1}{2},1)$. For every $t\geq 1$ denote
+$\delta_{s}(t):=\Re\big(\psi_{s}(1,t)-\psi_{1-\bar{s}}(1,t)\big).$
+Thank to the Proposition 4, we have $\dot{\delta}_{s}(1)>0$. By continuity there exists $\epsilon>0$ such that
+$\delta_{s}(t)>0,\quad\forall t\in(1,1+\epsilon).$
 Define the time $t_{*}>1$ as
-$t_{*}:=\sup\{t>1:\forall\ 1<u<t,\ x_{s}(u)>x_{s^{\prime}}(u)\}.$ (6)
-The strategy to prove the present Proposition is to prove that $t_{*}=+\infty$. By contradiction, suppose that $t_{*}<+\infty$. Then
-$x_{s}(t_{*})=x_{s^{\prime}}(t_{*}),$
+$t_{*}:=\sup\{t>1:\forall\ 1<u<t,\ \delta_{s}(u)>0\}.$
+To prove of the present Proposition it is equivalent to prove that $t_{*}=+\infty$. By contradiction, suppose that $t_{*}<+\infty$. Then
+$\delta_{s}(t_{*})=0,$ (6)
 and we have
-$x_{s}(t)>x_{s^{\prime}}(t),\quad\forall t\in(1,t_{*}).$ (7)
-We have two cases: $x_{s}(t_{*})\geq 0$ and $x_{s}(t_{*})<0$.
-If $x_{s}(t_{*})\geq 0$, then $x_{s^{\prime}}(t_{*})\geq 0$. Using the fact that $\sigma>\sigma^{\prime}$, from the differential equations (5), we get
-$\dot{x}_{s}(t_{*})-\dot{x}_{s^{\prime}}(t_{*})$ $=(\sigma-\sigma^{\prime})t_{*}^{-1}x_{s}(t_{*})$
-$+(\sigma-\sigma^{\prime})t_{*}^{-1}\big(1-\Re\big(\,t_{*}^{-i\tau}\,\eta(t_{*})\big)\big)$
-$\geq(\sigma-\sigma^{\prime})t_{*}^{-1}\big(1-\Re\big(\,t_{*}^{-i\tau}\,\eta(t_{*})\big)\big).$
-By the hypothesis (H), we obtain $\dot{x}_{s}(t_{*})>\dot{x}_{s^{\prime}}(t_{*})$.
-There exists $t<t_{*}$ close enough to $t_{*}$ such that $x_{s}(t)<x_{s^{\prime}}(t)$.
-We have obtained a contradiction with equation (7).
-If $x_{s}(t_{*})<0$, then $x_{s^{\prime}}(t_{*})<0$. By continuity there exists $\tilde{\epsilon}>0$ such that
-$0>x_{s}(u)>x_{s^{\prime}}(u),\quad\forall u\in(t_{*}-\tilde{\epsilon},t_{*}).$
+$\delta_{s}(u)>0,\quad\forall u\in(1,t_{*}).$ (7)
+Thank to the Proposition 4, we have
+${\delta}_{s}(t)=$ $\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
+$+4(\sigma-\frac{1}{2})\Re\!\left(\psi_{\frac{1}{2}+i\,\tau}(1,t)-\psi_{i\tau}(1,t)\right).$
+Since $2\sigma-1>0$, thanks to the hypothesis H, we obtain
+${\delta}_{s}(t)>\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du,\quad\forall t>1.$ (8)
+We obtain
+${\delta}_{s}(t_{*})>\frac{1}{2}\int_{1}^{t_{*}}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t_{*}}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du,$
+Since $2\sigma-1>0$, by equation (7) we get
+${\delta}_{s}(t_{*})>0$
+We obtained a contradiction with equation (6). Then $t_{*}=+\infty$. In other words, $\delta_{s}(t)>0$ for all $t>1$. The equation (8) implies
+$\dot{\delta}_{s}(t)>(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{2}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du,\quad\forall t>1.$
+That gives
+${\delta}_{s}(t)>2(\sigma-\frac{1}{2})^{2}(t^{\frac{1}{2}}-1)\int_{1}^{2}u^{-\frac{3}{2}}{\delta}_{s}(u)du,\quad\forall t>1.$
+Then $\lim_{t\to+\infty}\delta_{s}(t)=+\infty$.
+This implies that the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ are not both bounded on $1,+\infty)$.
+By Lemma 3, we deduce that
+$(1,1)\neq\big(\mu(s),\,\mu(1-\overline{s})\big).$
+More precisely, the proof of Proposition 4 shows that under the rigidity hypothesis H, for all $s\in B$ with $\Re(s)\in(\frac{1}{2},1)$, the function $\mu$ satisfies the following equation
+$\Re\big(\mu(s)\big)-\Re\big(\mu(1-\overline{s})\big)\leq 2\Re\big(\eta(1)\big).$
+In this case, $\dot{\delta}_{s}(1)\leq 0$; That is, the function does not necessary start positively from 11. Hence the proof of the Theorem 5 cannot be continued.
+3 Application to the $\zeta$-Function
+The function $\eta(t)=\{t\}$, the fractional part of $t$, is a locally integrable function and satisfies $0\leq\eta(t)<1$.
+We will now see that hypothesis H, is also satisfied.
+From the equation (2), we have
+$\psi_{w}(1,t)=t^{\Re(w)}\Big(1+(1-w)\int_{1}^{t}u^{-1-w}\,\{u\}\,du\Big),\quad\forall w\in\mathbb{C},\quad\forall t>1,$
+which can be written as
+$\psi_{w}(1,t)$ $=t^{\Re(c)}\Big[1+J(t,\lfloor t\rfloor)+\sum_{n=1}^{\lfloor t\rfloor-1}J(n+1,n)\Big],$
+$J_{w}(x,n)$ $:=(1-w)\int_{n}^{x}u^{-1-w}\,(u-n)\,du,$
+with $\sum_{n=1}^{\lfloor t\rfloor-1}J_{w}(n+1,n)=0$ when $t\in[1,2)$. Use the integration by parts formula, we obtain
+$J_{w}(x,n)$ $=\int_{n}^{x}\Big(x^{1-w}-u^{1-w}\Big)\,\frac{d}{du}\Big(u^{-1}\,(u-n)\Big)du$
+$=\int_{n}^{x}\Big(x^{1-w}-u^{1-w}\Big)\,u^{-2}du.$
+For every $\beta\in\mathbb{R}$ and $t\geq 2$ we have
+$\Big|t^{\frac{1}{2}}J_{\frac{1}{2}+i\,\beta}(x,n)-J_{i\,\beta}(x,n)\Big|$ $\leq\int_{n}^{x}\Big(\big(t^{\frac{1}{2}}x^{\frac{1}{2}}-x\big)+(t^{\frac{1}{2}}u^{\frac{1}{2}}-u)\,\Big)u^{-2}du$
 Implies
-$\delta_{s,s^{\prime}}(u):=x_{s}(u)x_{s^{\prime}}(u)-(x_{s}(u))^{2}=x_{s}(u)(x_{s^{\prime}}(u)-x_{s}(u))>0,\ \forall u\in(t_{*}-\tilde{\epsilon},t_{*}).$
-and $\delta_{s,s^{\prime}}(t_{*})=0$.
-Then we must find $\dot{\delta}_{s,s^{\prime}}(t_{*})\leq 0$.
-As the previous technique, prove that in our system that $\dot{\delta}_{s,s^{\prime}}(t_{*})>0$.
-From the differential equation (4), we have
-[…]
-$x_{s^{\prime}}(t)\dot{x}_{s}(t)$ $=\sigma\,t^{-1}x_{s^{\prime}}(t)x_{s}(t)+\sigma t^{-1}x_{s^{\prime}}(t)+t^{-1}x_{s^{\prime}}(t)\Re\big((1-s)\,t^{-i\tau}\,\eta(t)\big),$
-$x_{s}(t)\dot{x}_{s^{\prime}}(t)$ $=\sigma^{\prime}\,t^{-1}x_{s}(t)x_{s^{\prime}}(t)+\sigma^{\prime}t^{-1}x_{s}(t)+t^{-1}x_{s}(t)\Re\big((1-s^{\prime})\,t^{-i\tau}\,\eta(t)\big).$
-Implies
-$\dot{\delta}_{s,s^{\prime}}(t)$ $=2\sigma\,t^{-1}{\delta}_{s,s^{\prime}}(t)+t^{-1}\big(\sigma x_{s^{\prime}}(t)+\sigma^{\prime}x_{s}(t)\big)$
-$+t^{-1}x_{s^{\prime}}(t)\Re\big((1-s)\,t^{-i\tau}\,\eta(t)\big)$
-$-t^{-1}x_{s}(t)\Re\big((1-s^{\prime})\,t^{-i\tau}\,\eta(t)\big)$
-$-2\sigma t^{-1}x_{s}(t).$
-At the point $t_{*}$ we get
-$\dot{\delta}_{s,s^{\prime}}(t_{*})$ $=-t_{*}^{-1}\big(\sigma-\sigma^{\prime}\big)x_{s}(t_{*})\big(1-\Re\big(t_{*}^{-i\tau}\,\eta(t_{*})\big)\big).$
-Use the fact that $\sigma>\sigma^{\prime}$ and $x_{s}(t_{*})<0$, we obtain $\dot{\delta}_{s,s^{\prime}}(t_{*})>0$. Same, we obtained a contradiction. Then $t_{*}=+\infty$. ∎
-3 Proof of the Theorem 4
-Intuitively, the solution $\psi_{w}(1,\cdot)$cannot be bounded when $\Re(w)\in\left(\tfrac{1}{2},1\right)$.
-Indeed, if $\Re(w)\in(0,1)$, then |w|>|1w||w|>|1-w|.
-Since the initial condition satisfies $\psi_{w}(1,1)=1$ and $\sup_{t\geq 1}|\eta(t)|\leq 1$, for $t$ close to 11 the homogeneous part dominates the non-homogeneous part in differential equation (1).
-This may therefore lead to the unboundedness of the solution $\psi_{w}(1,\cdot)$.
-In this manuscript, we compare and study the boundedness of two solutions corresponding to two complex parameters with the same imaginary part, in order to eliminate the imaginary part (which may be very large) from the computations.
-Proof of the Theorem 4.
-We prove the theorem by contradiction.
-Suppose there exist $s,s^{\prime}\in B$ with $\Re(s)>\Re(s^{\prime})$ and $\Im(s)=\Im(s^{\prime})$ such that $\big(\mu(s),\,\mu(s^{\prime})\big)=(1,1)$.
-The solutions $\psi_{s}(1,\cdot)$ and $\psi_{s^{\prime}}(1,\cdot)$ of the differential equation (1) are both bounded on $1,+\infty)$.
-Thanks to the second Item of the Lemma 3, we have
-$\sup_{t\geq 1}\Big|t\Big(\psi_{s}(1,t)+\rho\frac{1-s}{s}\Big)\Big|<+\infty\ \text{and}\ \sup_{t\geq 1}\Big|t\Big(\psi_{s^{\prime}}(1,t)+\rho\frac{1-s^{\prime}}{s^{\prime}}\Big)\Big|<+\infty.$
-Use the Proposition 5, we get
-$\lim_{t\to+\infty}\Re\Big(t^{-i\Im(s)}\rho\big(\frac{1-s^{\prime}}{s^{\prime}}-\frac{1-s}{s}\big)\Big)\geq 0.$
-Since $s,s^{\prime}\in B$, then $\Im(s)\neq 0$, that gives
-$\rho\big(\frac{1-s^{\prime}}{s^{\prime}}-\frac{1-s}{s}\big)=0.$
-By the hypothesis (H), we have $\rho\neq 0$. then
-$\frac{1-s^{\prime}}{s^{\prime}}-\frac{1-s}{s}=0$
-We have obtained a contradiction with $\Re(s)\neq\Re(s^{\prime})$. ∎
-4 Application to the $\zeta$-Function
-The function $\eta(t)=\{t\}$ (the fractional part of $t$) is a locally integrable function and satisfies $0\leq\eta(t)<1$ with a rotation number $\rho$ (here $\rho=1/2$).
-The Hypothesis (H) is satisfied.
-By definition of the function $\mu$ in the Definition 2, we have
+$\psi_{\frac{1}{2}+i\,\beta}(1,t)-\psi_{i\,\beta}(1,t)$ $\geq t^{\frac{1}{2}}-1-\int_{\lfloor t\rfloor-1}^{t}(t^{\frac{1}{2}}u^{\frac{1}{2}}-u)u^{-2}du$
+$+\sum_{n=1}^{\lfloor t\rfloor}\int_{n}^{n+1}\big(n+1+u)\,u^{-2}du$
+$-t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}\int_{n}^{n+1}\big((n+1)^{\frac{1}{2}}+u^{\frac{1}{2}}\big)u^{-2}du$
+$\geq\frac{1}{3}\big(t^{\frac{1}{2}}-1\big)+\left(1-\frac{1}{t}\right),\qquad\forall t\geq 1.$
+We recall that $\sum_{n=1}^{\lfloor t\rfloor-1}J(n+1,n)=0$ for all $t\in[1,2)$. We deduce that
+$\Re\Big(\psi_{\frac{1}{2}+i\,\beta}(1,t)-\psi_{i\,\beta}(1,t)\Big)>0,\quad\forall t>1.$
+Then the Hypothesis hypothesis H is satisfied. By definition of the function $\mu$ in the Definition 2, we have
 $\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\{u\}\,du,\quad\mu(1-\bar{s})=-\bar{s}\int_{1}^{+\infty}u^{-2+\bar{s}}\{u\}\,du.$
@@
 $\frac{1-s}{s}\zeta(s):=-1+\mu(s),\quad\frac{\bar{s}}{1-\bar{s}}\zeta(1-\bar{s}):=-1+\mu(1-\bar{s}).$ (9)
-Thanks to the Riemann functional equation [1, Theorem 2.1, p.
-13], when $\zeta(s)=0$ then $\zeta(1-\bar{s})=0$.
-Implies $\big(\mu(s),\ \mu(1-\bar{s})\big)=(1,1)$.
-By the Theorem 4 we deduce that $\Re(s)=\Re(1-\bar{s})$.
+Since the Hypothesis hypothesis H is satisfied, by the Theorem 5 when $s\in B$ with $\Re(s)\neq\frac{1}{2}$ we get $\big(\mu(s),\ \mu(1-\bar{s})\big)\neq(1,1)$.
 The non-trivial zeros of the Riemann zeta function must lie on the critical line.

### v6

@@
-Structural Asymmetry and Rigidity of Bounded Solutions for a Parametrized Complex Linear Differential Equation
+Rigidity Hypothesis for the Structural Asymmetry of the Bounded Solutions of a Parametric Complex Linear Differential Equation
 Abstract
-In this manuscript, we introduce a family of parametrized nonhomogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter lying in the critical strip.
-We identify a suitable rigidity hypothesis under which a phenomenon of structural asymmetry is established: if two solutions with the same initial condition equal to 11, corresponding respectively to the parameters $s$ and $1-\bar{s}$, are both bounded on $1,+\infty)$, then $\Re(s)=\frac{1}{2}$.
+In this manuscript, we introduce a family of parametrized non-homogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter lying in the critical strip.
+We identify a rigidity hypothesis along the critical line, which establishes a structural asymmetry: if two solutions with the same initial condition equal to 11, corresponding respectively to the parameters $s$ and $1-\bar{s}$, are both bounded on $1,+\infty)$, then $\Re(s)=\frac{1}{2}$.
+This rigidity hypothesis is satisfied when the non-homogeneous term is the fractional part function.
 Keywords: Non-homogeneous complex linear differential equation, bounded solutions.
@@
 1 Main results
-In this manuscript, we study the following non-homogeneous linear complex parametrized differential equations
+In this manuscript, we study the following non-homogeneous linear complex differential equations:
+$\dot{\theta}=wt^{-1}\theta+t^{-1}\eta(t),\quad\theta:[1,+\infty)\to\mathbb{C},$
+where $w\in\mathbb{C}$ is the parameter and the non-homogeneous term $\eta$ is a normalized, locally integrable function satisfying $|\eta(t)|<1$ for all $t\geq 1$.
+We are interested in the initial conditions of the bounded solutions.
+To this end, we analyze the behavior of the transformed solution $t\mapsto(1-w)t^{-i\Im(w)}\theta(t)$, which leads us to consider the following differential equation:
 $\dot{x}=\Re(w)t^{-1}x+(1-w)t^{-1-i\Im(w)}\eta(t),$ (1)
-$t\in[1,+\infty),\quad x(1)=z,\quad x:[1,+\infty)\to\mathbb{C},$
-where $w\in\mathbb{C}$ and $\eta:[1,+\infty)\to\mathbb{C}$ is a locally integrable function and satisfy $|\eta(t)|<1$ for all $t\geq 1$.
-The equation (1) can be viewed as a differential equation defined on $1,+\infty)$ with values in $\mathbb{R}^{2}$, and we are interested in the initial conditions of the bounded solutions.
-This result provides an answer to the Dynamical Conjecture formulated in [2].
-In the study that follows, we fix a specific function $\eta$ which is locally integrable function and satisfy $|\eta(t)|<1$ for all $t\geq 1$.
-The function
-$t\mapsto t^{\Re(w)}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1,$
-is continuous on $1,+\infty)$.
+$t\in[1,+\infty),\quad x(1)=z,\quad x:[1,+\infty)\to\mathbb{C}.$
+In the following study, we fix a specific function $\eta$ as defined above. The function
+$t\mapsto t^{\Re(w)}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1.$
+is absolutely continuous on $1,+\infty)$.
 The differential equation (1) is a non-homogeneous linear differential equation.
@@
 Notation 2.
-Denote by $B\subset\mathbb{C}$ the following set
-$B:=\Big\{w\in\mathbb{C}:\quad\Re(w)>0\Big\}.$
-Let $\mu$ denote the function $\mu:B\mapsto\mathbb{C}$, defined as
-$\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\eta(u)\,du,\quad\forall s\in B.$
+Denote by $\mathbb{C}_{+}$ the right half-plane defined as
+$\mathbb{C}_{+}:=\Big\{w\in\mathbb{C}:\quad\Re(w)>0\Big\}.$
+Let $\mu$ denote the function $\mu:\mathbb{C}_{+}\mapsto\mathbb{C}$, defined as
+$\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\eta(u)\,du,\quad\forall s\in\mathbb{C}_{+}.$
+The function $\mu$ is defined for all $s\in\mathbb{C}_{+}$.
+Indeed, the integral is absolutely convergent, since for all $w\in\mathbb{C}_{+}$ we have $\Re(w)>0$.
 In the following lemma, we are interested in bounded solutions on $1,+\infty)$ and their initial condition.
 Lemma 3.
-Let $w\in B$ and $z\in\mathbb{C}$. Consider the continuous solution $\psi_{w}(z,\cdot)$ of the differential equation (1) as given in Notation 1. Then
+Let $w\in\mathbb{C}_{+}$ and $z\in\mathbb{C}$. Consider the continuous solution $\psi_{w}(z,\cdot)$ of the differential equation (1) as given in Notation 1. Then
 $\sup_{t\geq 1}|\psi_{w}(z,t)|<+\infty\quad\text{if and only if}\quad z=\mu(w),$
@@
 Proof.
-Let $w\in B$ be fixed.
+Let $w\in\mathbb{C}_{+}$ be fixed.
 There is a unique bounded solution; all other solutions diverge in norm.
@@
 $\Big|z+(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du\Big|=0,$
-which gives $z=\mu(w)$. To prove that $\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|<+\infty$, observe that by definition of $\mu(w)$,
+which gives $z=\mu(w)$. Now, prove that that $\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|<+\infty$. By definition of $\mu(w)$, we have
 $\mu(w)+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du=-(1-w)\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$
@@
 By Hypothesis we have $\sup_{t\geq 1}|\eta(t)|\leq 1$, we obtain
-$\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|\leq\frac{|1-w|}{\Re(w)}\sup_{t\geq 1}|\eta(t)|<\frac{|1-w|}{\Re(w)}.$
-From the perspective of the qualitative theory of dynamical systems, the differential equation (1) can be viewed as a differential equation with the parameter $\Re(w)$.
+$\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|\leq\frac{|1-w|}{\Re(w)}\sup_{t\geq 1}|\eta(t)|\leq\frac{|1-w|}{\Re(w)}.$
+The differential equation (1) can be viewed as a differential equation with the parameter $\Re(w)$.
 Intuitively, this introduces a order structure within the set of solutions with a same initial condition and parameterized by $\Re(w)$.
-We show this in the following proposition.
+We start to exhibit this order between solutions through the following proposition.
 Proposition 4.
-Let be $s:\sigma+i\tau\in B$.
-Consider the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) as given in Notation 1.
-The function $\delta_{s}$ defined as
+Consider the continuous solutions of the differential equation (1) as given in Notation 1. For every $s:=\sigma+i\,\tau\in\mathbb{C}_{+}$ with $\sigma\in(0,1)$, the function $\delta_{s}$ defined as
 $\delta_{s}(t):=\Re\big({\psi_{s}}(1,t)-{\psi}_{1-\bar{s}}(1,t)\big),\quad\forall t\geq 1.$
-satisfies the following two equalities:
-$\dot{\delta}_{s}(1)$ $=(2\sigma-1)\Big(1-\Re\big(\eta(1)\big)\Big),$
-${\delta}_{s}(t)$ $=\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
-$\ +4(\sigma-\frac{1}{2})\Re\!\left(\psi_{\frac{1}{2}+i\,\tau}(1,t)-\psi_{i\tau}(1,t)\right),\qquad\forall t\geq 1.$
-According to the notation 1 consider the following rigidity hypothesis H:
-  • H.
-    For every $\beta\in\mathbb{R}$ and every $t>1$,
-    $\Re\!\left(\psi_{\frac{1}{2}+i\beta}(1,t)-\psi_{i\beta}(1,t)\right)>0.$
-The previous hypothesis suggest the rigidity of the sign. It is suitable in order to apply the results of this dynamical study to the Riemann zeta function.
-The following theorem shows that, under the rigidity hypothesis H, if $\Re(s)\neq\tfrac{1}{2}$, then the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) cannot both be bounded on $1,+\infty)$.
+satisfies the following integral equation:
+$\delta_{s}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{s}(v)\,dv\,du$
+$+(2\sigma-1)t^{\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}}\Re\big(\psi_{\frac{1}{2}+i\,\tau}(1,u)\big)\,du$
+$+(2\sigma-1)t^{\frac{1}{2}}\Re\Big(\frac{1-t^{-\frac{1}{2}}\psi_{\frac{1}{2}+i\,\tau}(1,t)}{\frac{1}{2}-i\,\tau}\Big),\qquad\qquad\forall t\geq 1.$
+Denote by $L$ the critical ligne difened as
+$L:=\{w\in\mathbb{C},\quad\Re(w)=\frac{1}{2}\}.$
+Given the previous proposition, we consider the following rigidity hypothesis:
+$\forall w\in L,\ \forall t\geq 1:\quad\Re\Big(\frac{1-t^{-\frac{1}{2}}\psi_{w}(1,t)}{1-w}+\int_{1}^{t}u^{-\frac{3}{2}}\psi_{w}(1,u)\,du\Big)\geq 0.$ (H)
+The previous hypothesis represents a rigidity of the sign along the critical line. It is suitable in order to apply the results of this dynamical study to the Riemann zeta function.
+The following theorem shows that, under the rigidity hypothesis (H), if $\Re(s)\neq\tfrac{1}{2}$, then the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) cannot both be bounded on $1,+\infty)$.
 Theorem 5.
-Consider the function $\mu$ defined in Notation 2. Under hypothesis H, for every $s\in\mathbb{C}$ with $\Re(s)\in(0,1)/\{\frac{1}{2}\}$, we have
+Consider the function $\mu$ given by Notation 2. Under hypothesis (H), for every $s\in\mathbb{C}$ with $\Re(s)\in(0,1)/\{\frac{1}{2}\}$, we have
 $\big(\mu(s),\,\mu(1-\overline{s})\big)\neq(1,1).$
@@
 Proof of the Proposition 4.
-Let be $w\in B$ and denote $w:=\alpha+i\beta$. For every $t\geq 1$ denote
+Let be $w\in\mathbb{C}_{+}$ and denote $w:=\alpha+i\,\beta$. For every $t\geq 1$ denote
 $x_{w}(t):=t^{-\frac{1}{2}}\Re\big(\psi_{w}(1,t)\big).$ (4)
@@
 $\dot{x}_{w}(t)=\big(\alpha-\frac{1}{2}\big)\,t^{-1}x_{w}(t)+t^{-\frac{3}{2}}\Re\big((1-w)\,t^{-i\,\beta}\,\eta(t)\big),\quad x_{w}(1)=1.$ (5)
-Let be $s:=\sigma+i\tau\in B$ such that $\sigma\in(\frac{1}{2},1)$. From the previous equation, we have
+Let be $s:=\sigma+i\,\tau\in\mathbb{C}_{+}$ such that $\sigma\in(0,1)$. From the previous equation, we have
 $\dot{x_{s}}(t)+\dot{x}_{1-\bar{s}}(t)=(\sigma-\frac{1}{2})t^{-1}\big({x_{s}}(t)-{x}_{1-\bar{s}}(t)\big)+t^{-\frac{3}{2}}\Re\Big((1-2i\tau)t^{-i\,\tau}\,\eta(t)\Big).$
@@
 $+(\sigma-\frac{1}{2})t^{-1}\Re\Big((1-2i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\,du\Big).$
-Define
-$\delta_{s}(t):=\Re\big({\psi_{s}}(1,t)-{\psi}_{1-\bar{s}}(1,t)\big),\quad\forall t\geq 1.$
-By the notation (4), we have ${x_{s}}(t)-{x}_{1-\bar{s}}(t):=t^{-\frac{1}{2}}\delta_{s}(t)$. We obtained
-$\dot{\delta}_{s}(t)=$ $\frac{1}{2}t^{-1}\delta_{s}(t)+(\sigma-\frac{1}{2})^{2}t^{-\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}}\delta_{s}(u)\,du$
-$+2(\sigma-\frac{1}{2})t^{-\frac{1}{2}}\Big(1-t^{-\frac{1}{2}}\Re\big(\,t^{-i\tau}\,\eta(t)\big)\Big)$
-$+2(\sigma-\frac{1}{2})t^{-\frac{1}{2}}\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\tau}\,\eta(u)\,du\Big).$
-Since ${\delta}_{s}(1)=0$ and since by hypothesis we have $|\eta(1)|<1$, then
-$\dot{\delta}_{s}(1)=(2\sigma-1)\Big(1-\Re\big(\eta(1)\big)\Big).$
-Now, prove the second item of the present Proposition. Integrate the previous equation, we obtain
-${\delta}_{s}(t)=\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}\dot{\delta}_{s}(v)\,dv\,du$
-$+2(\sigma-\frac{1}{2})\int_{1}^{t}u^{-\frac{1}{2}}\Re\Big(1-u^{-\frac{1}{2}-i\tau}\,\eta(u)\Big)du$
-$+2(\sigma-\frac{1}{2})\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}-i\tau}\,\eta(v)\,dv\,du\Big).$
-Use the integration by pars formula to the last integral, we get
-${\delta}_{s}(t)=\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}\dot{\delta}_{s}(v)\,dv\,du$
-$+2(\sigma-\frac{1}{2})\int_{1}^{t}u^{-\frac{1}{2}}\Re\Big(1-u^{-\frac{1}{2}-i\tau}\,\eta(u)\Big)du$
-$+4(\sigma-\frac{1}{2})t^{\frac{1}{2}}\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}v^{-\frac{3}{2}-i\tau}\,\eta(v)\,dv\Big)$
-$-4(\sigma-\frac{1}{2})\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-1-i\tau}\,\eta(u)\,du\Big).$
-Equivalent to
-${\delta}_{s}(t)=$ $\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
-$+4(\sigma-\frac{1}{2})t^{\frac{1}{2}}\big[(\frac{1}{2}-i\tau)\int_{1}^{t}v^{-\frac{3}{2}-i\tau}\,\eta(v)\,dv\Big]$
-$-4(\sigma-\frac{1}{2})\Big[1+\Re\Big((1-i\tau)\int_{1}^{t}u^{-1-i\tau}\,\eta(u)\,du\Big)\Big].$
-According to Notation 1, we obtain
-${\delta}_{s}(t)=$ $\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
-$+4(\sigma-\frac{1}{2})\Re\!\left(\psi_{\frac{1}{2}+i\beta}(1,t)-\psi_{i\beta}(1,t)\right).$
+Use the fact that ${x}_{s}(1)-{x}_{1-\bar{s}}(1)=0$ and integrate, we obtain
+${x}_{s}(t)-{x}_{1-\bar{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-1}\big({x_{s}}(v)-{x}_{1-\bar{s}}(v)\big)\,dv\,du$
+$+2(\sigma-\frac{1}{2})\Re\big(\,\int_{1}^{t}u^{-1}\Big(1-u^{-\frac{1}{2}}u^{-i\,\tau}\,\eta(u)\big)du\Big)$
+$+2(\sigma-\frac{1}{2})\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\,du\Big).$
+From the equation (2), for every $\beta\in\mathbb{R}$ and $t\geq 1$, we have
+$\psi_{\frac{1}{2}+i\,\beta}(1,t)=t^{\frac{1}{2}}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\eta(u)\,du\Big],\quad\forall\beta\in\mathbb{R}.$
+Therefore,
+${x}_{s}(t)-{x}_{1-\bar{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-1}\big({x_{s}}(v)-{x}_{1-\bar{s}}(v)\big)\,dv\,du$
+$+2(\sigma-\frac{1}{2})\int_{1}^{t}u^{-\frac{3}{2}}\Re\Big(\psi_{\frac{1}{2}+i\,\tau}(1,u)-u^{-i\,\tau}\,\eta(t)\Big)\,du,$
+and that can be written as
+${x}_{s}(t)-{x}_{1-\bar{s}}(t)$ $=(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-1}\big({x_{s}}(v)-{x}_{1-\bar{s}}(v)\big)\,dv\,du$
+$+(2\sigma-1)\int_{1}^{t}u^{-\frac{3}{2}}\Re\big(\psi_{\frac{1}{2}+i\,\tau}(1,u)\big)\,du$
+$+(2\sigma-1)\Re\Big(\frac{1-t^{-\frac{1}{2}}\psi_{\frac{1}{2}+i\,\tau}(1,t)}{\frac{1}{2}-i\,\tau}\Big).$
+By the notation (4), we have
+${x_{s}}(t)-{x}_{1-\bar{s}}(t)=t^{-\frac{1}{2}}\delta_{s}(t):=t^{-\frac{1}{2}}\Re\big({\psi_{s}}(1,t)-{\psi}_{1-\bar{s}}(1,t)\big).$
 Proof of the Theorem 5.
-Let be $s:=\sigma+i\tau\in B$ such that $\sigma\in(0,1)/\{\frac{1}{2}\}$. By symmetry, we may assume that $\sigma\in(\frac{1}{2},1)$. For every $t\geq 1$ denote
+Let be $s:=\sigma+i\,\tau\in\mathbb{C}$ such that $\sigma\in(0,1)/\{\frac{1}{2}\}$. By symmetry, we may assume that $\sigma\in(\frac{1}{2},1)$. For every $t\geq 1$ denote
 $\delta_{s}(t):=\Re\big(\psi_{s}(1,t)-\psi_{1-\bar{s}}(1,t)\big).$
-Thank to the Proposition 4, we have $\dot{\delta}_{s}(1)>0$. By continuity there exists $\epsilon>0$ such that
+From the equation (1) , we have
+$\dot{\delta}_{s}(t)=t^{-1}\Re\Big(\sigma\psi_{s}(1,t)-(1-\sigma)\psi_{1-\bar{s}}(1,t)\Big)-(2\sigma-1)\,t^{-1}\Re\big(t^{-i\,\tau}\eta(t)\Big),$
+Since $\psi_{s}(1,1)=\psi_{1-\bar{s}}(1,1)=1$ then ${\delta}_{s}(1)=0$.
+By hypothesis, we have $|\eta(t)|<1$ for all $t\geq 1$.
+Since $\sigma\in(\frac{1}{2},1)$ the previous differential equation implies $\dot{\delta}_{s}(1)>0$.
+By continuity there exists $\epsilon>0$ such that
 $\delta_{s}(t)>0,\quad\forall t\in(1,1+\epsilon).$
@@
 Thank to the Proposition 4, we have
-${\delta}_{s}(t)=$ $\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
-$+4(\sigma-\frac{1}{2})\Re\!\left(\psi_{\frac{1}{2}+i\,\tau}(1,t)-\psi_{i\tau}(1,t)\right).$
-Since $2\sigma-1>0$, thanks to the hypothesis H, we obtain
-${\delta}_{s}(t)>\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du,\quad\forall t>1.$ (8)
+$\delta_{s}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{s}(v)\,dv\,du$
+$+(2\sigma-1)t^{\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}}\Re\big(\psi_{\frac{1}{2}+i\,\tau}(1,u)\big)\,du$
+$+(2\sigma-1)t^{\frac{1}{2}}\Re\Big(\frac{1-t^{-\frac{1}{2}}\psi_{\frac{1}{2}+i\,\tau}(1,t)}{\frac{1}{2}-i\,\tau}\Big),\qquad\qquad\forall t\geq 1.$
+Since $2\sigma-1>0$, thanks to the hypothesis (H), we obtain
+${\delta}_{s}(t)\geq(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{s}(v)\,dv\,du,\quad\forall t\geq 1.$ (8)
 We obtain
-${\delta}_{s}(t_{*})>\frac{1}{2}\int_{1}^{t_{*}}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t_{*}}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du,$
+${\delta}_{s}(t_{*})\geq(\sigma-\frac{1}{2})^{2}t_{*}^{\frac{1}{2}}\int_{1}^{t_{*}}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{s}(v)\,dv\,du.$
 Since $2\sigma-1>0$, by equation (7) we get
-${\delta}_{s}(t_{*})>0$
+${\delta}_{s}(t_{*})>0,$
 We obtained a contradiction with equation (6). Then $t_{*}=+\infty$. In other words, $\delta_{s}(t)>0$ for all $t>1$. The equation (8) implies
-$\dot{\delta}_{s}(t)>(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{2}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du,\quad\forall t>1.$
-That gives
-${\delta}_{s}(t)>2(\sigma-\frac{1}{2})^{2}(t^{\frac{1}{2}}-1)\int_{1}^{2}u^{-\frac{3}{2}}{\delta}_{s}(u)du,\quad\forall t>1.$
+${\delta}_{s}(t)>(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{2}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{s}(v)\,dv\,du,\quad\forall t>1.$
 Then $\lim_{t\to+\infty}\delta_{s}(t)=+\infty$.
@@
 $(1,1)\neq\big(\mu(s),\,\mu(1-\overline{s})\big).$
-More precisely, the proof of Proposition 4 shows that under the rigidity hypothesis H, for all $s\in B$ with $\Re(s)\in(\frac{1}{2},1)$, the function $\mu$ satisfies the following equation
-$\Re\big(\mu(s)\big)-\Re\big(\mu(1-\overline{s})\big)\leq 2\Re\big(\eta(1)\big).$
-In this case, $\dot{\delta}_{s}(1)\leq 0$; That is, the function does not necessary start positively from 11. Hence the proof of the Theorem 5 cannot be continued.
+Remark 6.
+  • Under hypothesis (H), the process used in the proofs of Proposition 4 and Theorem 5 shows that: For all $s\in\mathbb{C}$ with $\Re(s)\in(\frac{1}{2},1)$, the function $\mu$ satisfies the following implication
+    $\Re\Big(\mu(s)+\mu(1-\overline{s})\Big)\geq 2\implies\Re\Big(\mu(s)-\mu(1-\overline{s})\Big)<0.$
+    In this case, the function ${\delta}_{s}$ start with a negative values at 11. Hence, the proof of Theorem 5 cannot be continued.
+  • We recall that the rigidity hypothesis (H) is given as:
+    $\forall w\in L,\ \forall t\geq 1:\quad\Re\Big(\frac{1-t^{-\frac{1}{2}}\psi_{w}(1,t)}{1-w}+\int_{1}^{t}u^{-\frac{3}{2}}\psi_{w}(1,u)\,du\Big)\geq 0.$
+    From the equation (2), for every $\beta\in\mathbb{R}$ and $t\geq 1$, we have
+    $\psi_{\frac{1}{2}+i\,\beta}(1,t)=t^{\frac{1}{2}}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\eta(u)\,du\Big],$
+    Hence, the rigidity hypothesis (H) is equivalent to:
+    $\forall\beta\in\mathbb{R},\quad\forall t\geq 1:\quad\Re\big(f_{\beta}(t)\big)\geq 0,$
+    where
+    $f_{\beta}(t):=$ $\ln(t)-\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\eta(u)\,du$
+    $+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\eta(v)\,dv\,du.$
 3 Application to the $\zeta$-Function
 The function $\eta(t)=\{t\}$, the fractional part of $t$, is a locally integrable function and satisfies $0\leq\eta(t)<1$.
-We will now see that hypothesis H, is also satisfied.
-From the equation (2), we have
-$\psi_{w}(1,t)=t^{\Re(w)}\Big(1+(1-w)\int_{1}^{t}u^{-1-w}\,\{u\}\,du\Big),\quad\forall w\in\mathbb{C},\quad\forall t>1,$
+We will see in the following proposition that the hypothesis (H) is also satisfied
+Proposition 7.
+When $\eta(t)=\{t\}$ the rigidity hypothesis (H) is satisfied.
+Proof.
+Given the second item of Remark 6, we prove that : For all $\beta\in\mathbb{R}$ and $t\geq 1$ we have $\Re\big(f_{\beta}(t)\big)\geq 0$, where
+$f_{\beta}(t):=\ln(t)-\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\,du.$
+We have
+$f_{\beta}(t)=-\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\Big]\,du,$
+equivalent to
+$f_{\beta}(t)=\int_{1}^{t}u^{-\frac{1}{2}}h_{\beta}(u)\,du=t^{-\frac{1}{2}}\int_{1}^{t}h_{\beta}(u)\,du+\frac{1}{2}\int_{1}^{t}u^{-\frac{3}{2}}\int_{1}^{u}h_{\beta}(v)\,dv\,du,$
+where
+$h_{\beta}(u):=-u^{-1-i\,\beta}\,\{u\}+u^{-\frac{1}{2}}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\Big].$
+Then it is sufficient to prove that $\int_{1}^{t}\Re\big(h_{\beta}(u)\big)\,du\geq 0$ for all $t\geq 1$. We have
+$\int_{1}^{t}h_{\beta}(u)\,du$ $=2t^{\frac{1}{2}}-2-\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du$
+$+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\,du$
+$=2t^{\frac{1}{2}}-2-\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du$
+$+2(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du-2(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du,$
+We obtained
+$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1+(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du$
+$-(1-i\beta)\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du.$
+Since $\{u\}=u-\lfloor u\rfloor$, and since
+$(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{1}{2}-i\,\beta}\,du-(1-i\beta)\int_{1}^{t}u^{-i\,\beta}\,\,du=0,$
+then
+$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\lfloor u\rfloor\,du$
+$+(1-i\beta)\int_{1}^{t}u^{-1-i\,\beta}\,\lfloor u\rfloor\,du,$
 which can be written as
-$\psi_{w}(1,t)$ $=t^{\Re(c)}\Big[1+J(t,\lfloor t\rfloor)+\sum_{n=1}^{\lfloor t\rfloor-1}J(n+1,n)\Big],$
-$J_{w}(x,n)$ $:=(1-w)\int_{n}^{x}u^{-1-w}\,(u-n)\,du,$
-with $\sum_{n=1}^{\lfloor t\rfloor-1}J_{w}(n+1,n)=0$ when $t\in[1,2)$. Use the integration by parts formula, we obtain
-$J_{w}(x,n)$ $=\int_{n}^{x}\Big(x^{1-w}-u^{1-w}\Big)\,\frac{d}{du}\Big(u^{-1}\,(u-n)\Big)du$
-$=\int_{n}^{x}\Big(x^{1-w}-u^{1-w}\Big)\,u^{-2}du.$
-For every $\beta\in\mathbb{R}$ and $t\geq 2$ we have
-$\Big|t^{\frac{1}{2}}J_{\frac{1}{2}+i\,\beta}(x,n)-J_{i\,\beta}(x,n)\Big|$ $\leq\int_{n}^{x}\Big(\big(t^{\frac{1}{2}}x^{\frac{1}{2}}-x\big)+(t^{\frac{1}{2}}u^{\frac{1}{2}}-u)\,\Big)u^{-2}du$
-Implies
-$\psi_{\frac{1}{2}+i\,\beta}(1,t)-\psi_{i\,\beta}(1,t)$ $\geq t^{\frac{1}{2}}-1-\int_{\lfloor t\rfloor-1}^{t}(t^{\frac{1}{2}}u^{\frac{1}{2}}-u)u^{-2}du$
-$+\sum_{n=1}^{\lfloor t\rfloor}\int_{n}^{n+1}\big(n+1+u)\,u^{-2}du$
-$-t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}\int_{n}^{n+1}\big((n+1)^{\frac{1}{2}}+u^{\frac{1}{2}}\big)u^{-2}du$
-$\geq\frac{1}{3}\big(t^{\frac{1}{2}}-1\big)+\left(1-\frac{1}{t}\right),\qquad\forall t\geq 1.$
-We recall that $\sum_{n=1}^{\lfloor t\rfloor-1}J(n+1,n)=0$ for all $t\in[1,2)$. We deduce that
-$\Re\Big(\psi_{\frac{1}{2}+i\,\beta}(1,t)-\psi_{i\,\beta}(1,t)\Big)>0,\quad\forall t>1.$
-Then the Hypothesis hypothesis H is satisfied. By definition of the function $\mu$ in the Definition 2, we have
+$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1$
+$-(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\Big[\lfloor t\rfloor\int_{\lfloor t\rfloor}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\,du+\sum_{n=1}^{\lfloor t\rfloor-1}n\int_{n}^{n+1}\,u^{-\frac{3}{2}-i\,\beta}\,du\Big]$
+$+(1-i\beta)\Big[\lfloor t\rfloor\int_{\lfloor t\rfloor}^{t}u^{-1-i\,\beta}\ du+\sum_{n=1}^{\lfloor t\rfloor-1}n\int_{n}^{n+1}\,u^{-1-i\,\beta}\,du\Big].$
+Integrate, to obtain
+$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-t^{\frac{1}{2}}\frac{\frac{1}{2}-i\beta}{-\frac{1}{2}-i\,\beta}\lfloor t\rfloor\big(t^{-\frac{1}{2}-i\,\beta}-\lfloor t\rfloor^{-\frac{1}{2}-i\,\beta}\big)$
+$+\frac{1-i\beta}{-i\,\beta}\lfloor t\rfloor\big(t^{-i\,\beta}-\lfloor t\rfloor^{-i\,\beta}\big)$
+$-\frac{\frac{1}{2}-i\beta}{-\frac{1}{2}-i\,\beta}t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor-1}n\big((n+1)^{-\frac{1}{2}-i\,\beta}-n^{-\frac{1}{2}-i\,\beta}\big)$
+$+\frac{1-i\beta}{-i\,\beta}\sum_{n=1}^{\lfloor t\rfloor-1}n\big((n+1)^{-i\,\beta}-n^{-i\,\beta}\big).$
+Using telescopic series, one obtains
+$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1+\Big(\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}-\frac{1-i\beta}{i\,\beta}\Big)\lfloor t\rfloor t^{-i\,\beta}$
+$-\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}\Big(1+\sum_{n=1}^{\lfloor t\rfloor-1}(n+1)^{-\frac{1}{2}-i\,\beta}\Big)$
+$+\frac{1-i\beta}{i\,\beta}\Big(1+\sum_{n=1}^{\lfloor t\rfloor-1}(n+1)^{-i\,\beta}\Big),$
+which is equivalent to,
+$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1+\Big(\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}-\frac{1-i\beta}{i\,\beta}\Big)\lfloor t\rfloor t^{-i\,\beta}$
+$-\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}+\frac{1-i\beta}{i\,\beta}\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}.$
+or,
+$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-\Big(\frac{1}{\frac{1}{2}+i\,\beta}-\frac{1}{i\,\beta}\Big)\{t\}t^{-i\,\beta}$
+$+t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}-\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}$
+$+\Big(\frac{1}{\frac{1}{2}+i\,\beta}-\frac{1}{i\,\beta}\Big)t^{1-i\,\beta}$
+$-\frac{1}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}+\frac{1}{i\,\beta}\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}.$
+Use the Abel summation formula to the last line, we obtain
+$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-\Big(\frac{1}{\frac{1}{2}+i\,\beta}-\frac{1}{i\,\beta}\Big)\{t\}t^{-i\,\beta}$
+$+t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}-\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}$
+$+\frac{1}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}+t^{\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}-i\beta}\{u\}\,du$
+$-\frac{1}{1-i\,\beta}-\int_{1}^{t}u^{-1-i\beta}\{u\}\,du.$
+A direct estimate show that $\Re\big(f_{\beta}(t)\big)>0$, for all $t>1$. ∎
+Then the rigidity hypothesis (H) is satisfied. By definition of the function $\mu$ in the Definition 2, we have
 $\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\{u\}\,du,\quad\mu(1-\bar{s})=-\bar{s}\int_{1}^{+\infty}u^{-2+\bar{s}}\{u\}\,du.$
@@
 $\frac{1-s}{s}\zeta(s):=-1+\mu(s),\quad\frac{\bar{s}}{1-\bar{s}}\zeta(1-\bar{s}):=-1+\mu(1-\bar{s}).$ (9)
-Since the Hypothesis hypothesis H is satisfied, by the Theorem 5 when $s\in B$ with $\Re(s)\neq\frac{1}{2}$ we get $\big(\mu(s),\ \mu(1-\bar{s})\big)\neq(1,1)$.
+Since the hypothesis (H) is satisfied, by the Theorem 5 when $s\in B$ with $\Re(s)\neq\frac{1}{2}$ we get $\big(\mu(s),\ \mu(1-\bar{s})\big)\neq(1,1)$.
 The non-trivial zeros of the Riemann zeta function must lie on the critical line.
+The results provides an answer to the Dynamical Conjecture formulated in [2].

### v7

@@
 We identify a rigidity hypothesis along the critical line, which establishes a structural asymmetry: if two solutions with the same initial condition equal to 11, corresponding respectively to the parameters $s$ and $1-\bar{s}$, are both bounded on $1,+\infty)$, then $\Re(s)=\frac{1}{2}$.
-This rigidity hypothesis is satisfied when the non-homogeneous term is the fractional part function.
+This provides a theoretical framework when the non-homogeneous term is the fractional part function.
 Keywords: Non-homogeneous complex linear differential equation, bounded solutions.
@@
 The function $\eta(t)=\{t\}$, the fractional part of $t$, is a locally integrable function and satisfies $0\leq\eta(t)<1$.
-We will see in the following proposition that the hypothesis (H) is also satisfied
-Proposition 7.
-When $\eta(t)=\{t\}$ the rigidity hypothesis (H) is satisfied.
-Proof.
-Given the second item of Remark 6, we prove that : For all $\beta\in\mathbb{R}$ and $t\geq 1$ we have $\Re\big(f_{\beta}(t)\big)\geq 0$, where
-$f_{\beta}(t):=\ln(t)-\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\,du.$
-We have
-$f_{\beta}(t)=-\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\Big]\,du,$
-equivalent to
-$f_{\beta}(t)=\int_{1}^{t}u^{-\frac{1}{2}}h_{\beta}(u)\,du=t^{-\frac{1}{2}}\int_{1}^{t}h_{\beta}(u)\,du+\frac{1}{2}\int_{1}^{t}u^{-\frac{3}{2}}\int_{1}^{u}h_{\beta}(v)\,dv\,du,$
-where
-$h_{\beta}(u):=-u^{-1-i\,\beta}\,\{u\}+u^{-\frac{1}{2}}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\Big].$
-Then it is sufficient to prove that $\int_{1}^{t}\Re\big(h_{\beta}(u)\big)\,du\geq 0$ for all $t\geq 1$. We have
-$\int_{1}^{t}h_{\beta}(u)\,du$ $=2t^{\frac{1}{2}}-2-\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du$
-$+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\,du$
-$=2t^{\frac{1}{2}}-2-\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du$
-$+2(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du-2(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du,$
-We obtained
-$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1+(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du$
-$-(1-i\beta)\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du.$
-Since $\{u\}=u-\lfloor u\rfloor$, and since
-$(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{1}{2}-i\,\beta}\,du-(1-i\beta)\int_{1}^{t}u^{-i\,\beta}\,\,du=0,$
-then
-$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\lfloor u\rfloor\,du$
-$+(1-i\beta)\int_{1}^{t}u^{-1-i\,\beta}\,\lfloor u\rfloor\,du,$
-which can be written as
-$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1$
-$-(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\Big[\lfloor t\rfloor\int_{\lfloor t\rfloor}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\,du+\sum_{n=1}^{\lfloor t\rfloor-1}n\int_{n}^{n+1}\,u^{-\frac{3}{2}-i\,\beta}\,du\Big]$
-$+(1-i\beta)\Big[\lfloor t\rfloor\int_{\lfloor t\rfloor}^{t}u^{-1-i\,\beta}\ du+\sum_{n=1}^{\lfloor t\rfloor-1}n\int_{n}^{n+1}\,u^{-1-i\,\beta}\,du\Big].$
-Integrate, to obtain
-$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-t^{\frac{1}{2}}\frac{\frac{1}{2}-i\beta}{-\frac{1}{2}-i\,\beta}\lfloor t\rfloor\big(t^{-\frac{1}{2}-i\,\beta}-\lfloor t\rfloor^{-\frac{1}{2}-i\,\beta}\big)$
-$+\frac{1-i\beta}{-i\,\beta}\lfloor t\rfloor\big(t^{-i\,\beta}-\lfloor t\rfloor^{-i\,\beta}\big)$
-$-\frac{\frac{1}{2}-i\beta}{-\frac{1}{2}-i\,\beta}t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor-1}n\big((n+1)^{-\frac{1}{2}-i\,\beta}-n^{-\frac{1}{2}-i\,\beta}\big)$
-$+\frac{1-i\beta}{-i\,\beta}\sum_{n=1}^{\lfloor t\rfloor-1}n\big((n+1)^{-i\,\beta}-n^{-i\,\beta}\big).$
-Using telescopic series, one obtains
-$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1+\Big(\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}-\frac{1-i\beta}{i\,\beta}\Big)\lfloor t\rfloor t^{-i\,\beta}$
-$-\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}\Big(1+\sum_{n=1}^{\lfloor t\rfloor-1}(n+1)^{-\frac{1}{2}-i\,\beta}\Big)$
-$+\frac{1-i\beta}{i\,\beta}\Big(1+\sum_{n=1}^{\lfloor t\rfloor-1}(n+1)^{-i\,\beta}\Big),$
-which is equivalent to,
-$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1+\Big(\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}-\frac{1-i\beta}{i\,\beta}\Big)\lfloor t\rfloor t^{-i\,\beta}$
-$-\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}+\frac{1-i\beta}{i\,\beta}\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}.$
-or,
-$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-\Big(\frac{1}{\frac{1}{2}+i\,\beta}-\frac{1}{i\,\beta}\Big)\{t\}t^{-i\,\beta}$
-$+t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}-\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}$
-$+\Big(\frac{1}{\frac{1}{2}+i\,\beta}-\frac{1}{i\,\beta}\Big)t^{1-i\,\beta}$
-$-\frac{1}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}+\frac{1}{i\,\beta}\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}.$
-Use the Abel summation formula to the last line, we obtain
-$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-\Big(\frac{1}{\frac{1}{2}+i\,\beta}-\frac{1}{i\,\beta}\Big)\{t\}t^{-i\,\beta}$
-$+t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}-\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}$
-$+\frac{1}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}+t^{\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}-i\beta}\{u\}\,du$
-$-\frac{1}{1-i\,\beta}-\int_{1}^{t}u^{-1-i\beta}\{u\}\,du.$
-A direct estimate show that $\Re\big(f_{\beta}(t)\big)>0$, for all $t>1$. ∎
-Then the rigidity hypothesis (H) is satisfied. By definition of the function $\mu$ in the Definition 2, we have
+The following arguments suggests that the rigidity hypothesis (H) is satisfied.
+The exact calculation is framed within the context of number theory and the study of the discrete zeta function along the critical line.
+1. Asymptotic Invariance under Critical Oscillation ($\beta\to+\infty$):
+By a application of the Riemann-Lebesgue Lemma, the oscillatory integral terms vanish relative to the logarithmic term as $\beta$ increases. It follows that:
+$\lim_{\beta\to+\infty}\mathfrak{R}(f_{\beta}(t))=\ln(t).$
+Therefore, for each fixed $t>1$, the rigidity hypothesis (H) is asymptotically satisfied when $\beta\to+\infty$.
+2. Dominance of the Zeta Rigidity Term: The function $f_{\beta}$ can be written as
+$f_{\beta}(t)$ $=\ln(t)\Big(1+\big(\frac{1}{2}-i\beta\big)\int_{1}^{+\infty}u^{-\frac{3}{2}-i\beta}\{u\}du\Big)$
+$-\big(\frac{1}{2}-i\beta\big)\ln(t)\int_{t}^{+\infty}u^{-\frac{3}{2}-i\beta}\{u\}du$
+$-\int_{1}^{t}u^{-\frac{3}{2}-i\beta}\{u\}du-\big(\frac{1}{2}-i\beta\big)\int_{1}^{t}\ln(u)u^{-\frac{3}{2}-i\beta}\{u\}\,du.$
+Considering the representation of the Riemann zeta function $\zeta$ defined by Abel’s summation formula [[1], page 14, Equation 2.1.5], we have
+$\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\beta}\zeta(\frac{1}{2}+i\beta)=1+\big(\frac{1}{2}-i\beta\big)\int_{1}^{+\infty}u^{-\frac{3}{2}-i\beta}\{u\}du,$
+Hence, for large $t$, the dynamics of $\Re\big(f_{\beta}(t)\big)$ are governed by
+$\Re\Big(\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\beta}\zeta(\frac{1}{2}+i\beta)\Big)\ln(t)$
+Numerical investigation of the mapping $g(\beta)=-\Re\Big(\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\beta}\zeta(\frac{1}{2}+i\beta)\Big)$ reveals two fundamental properties:
+  • $g(\beta)\geq 0$ for all $\beta\in\mathbb{R}$.
+  • $g(\beta)=0$ if and only if $s=\frac{1}{2}+i\beta$ is a non-trivial zero of the Riemann zeta function.
+By definition of the function $\mu$ in the Definition 2, we have
 $\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\{u\}\,du,\quad\mu(1-\bar{s})=-\bar{s}\int_{1}^{+\infty}u^{-2+\bar{s}}\{u\}\,du.$
 In other words, by considering the integral representation of the Riemann zeta function $\zeta$ defined by Abel’s summation formula [[1], page 14, Equation 2.1.5], we have the relations
-$\frac{1-s}{s}\zeta(s):=-1+\mu(s),\quad\frac{\bar{s}}{1-\bar{s}}\zeta(1-\bar{s}):=-1+\mu(1-\bar{s}).$ (9)
-Since the hypothesis (H) is satisfied, by the Theorem 5 when $s\in B$ with $\Re(s)\neq\frac{1}{2}$ we get $\big(\mu(s),\ \mu(1-\bar{s})\big)\neq(1,1)$.
-The non-trivial zeros of the Riemann zeta function must lie on the critical line.
-The results provides an answer to the Dynamical Conjecture formulated in [2].
+$\frac{1-s}{s}\zeta(s):=-1+\mu(s),\quad\frac{\bar{s}}{1-\bar{s}}\zeta(1-\bar{s}):=-1+\mu(1-\bar{s}).$
+The hypothesis (H) and the Theorem 5 implies that when $s\in B$ with $\Re(s)\neq\frac{1}{2}$ we get: $\big(\mu(s),\ \mu(1-\bar{s})\big)\neq(1,1)$.
+This suggests that the non-trivial zeros of the Riemann zeta function must lie on the critical line, providing an answer to the Dynamical Conjecture formulated in [2].